OpenStudy (anonymous):

Fan and medal The side of a square is 2^72. Using the area formula A = s^2, determine the area of the square.

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

s = (2^72) A = s^2 = ?

OpenStudy (anonymous):

sry 2 ^7/2

OpenStudy (anonymous):



Vocaloid (vocaloid):

well, the same idea applies... s = 2^(7/2) A = s^2 = ?

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

so 7 * 2 is 14

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

if s = 2^(7/2) then s squared = 2^(7/2) * 2^(7/2) = ?